OpenStudy (anonymous):

How do i solve Sin(2x)^2-Cos(2x)=Cos(2x) ... so far i've gotten to the step of 4(1-cos^x)(cos^2x)=2Cos(2x) what do i do next?

OpenStudy (anonymous):

\(sin(2x)^2=2\cos 2x\) \(sin(2x)^2-2\cos 2x=0\) \(1-\cos(2x)^2-2\cos(2x)=0\) \(-\cos(2x)^2-2\cos(2x)+1=0\) Now, let \(cos(2x)=t\). Then: \(-t^2-2t+1=0\) solve this for \(t\) you will get: \(t=\huge \frac{-2\mp2\sqrt{2}}{2}\) now make back to x. and solve for x for each of t values

OpenStudy (anonymous):

@myko Thank you! i will work it using your method and see what what i get :) Thank you very much!